OpenStudy (anonymous):

F(x)=3x+4 how can i graph that with the inverse.....

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

from chapter 12

OpenStudy (immanuelv):

f^-1 (x) = (x-4 )/ (3)

OpenStudy (anonymous):

\[f \left( x \right)=3x+4\]

OpenStudy (anonymous):

how will i graph that

OpenStudy (anonymous):

look at it like y=mx+b or y=3x+4

OpenStudy (anonymous):

i kno...but like can you hlep me find the Points i can Use to graph

OpenStudy (anonymous):

put your pincle on the y intercept of 4 move up 3 boxes and to the right 1

OpenStudy (anonymous):

its a chapter 12 problem i need to graph tht and the inverse

OpenStudy (immanuelv):

i told u the inverse

OpenStudy (anonymous):

i kno....but i need to get the Normal graph Correct....

OpenStudy (immanuelv):

it is \[1±(\sqrt(3-x))/(\sqrt(2))\]

OpenStudy (anonymous):

is the square root on the 3 only???

OpenStudy (immanuelv):

it is the answer for f(x)= -2(x+1)^2 +3

OpenStudy (immanuelv):

and the square root is for the whole thing!

OpenStudy (anonymous):

ok

OpenStudy (immanuelv):

square root of 3-x

OpenStudy (anonymous):

kk

OpenStudy (immanuelv):

:)

OpenStudy (immanuelv):

hope it helped!

OpenStudy (anonymous):

it did but i need to show my work...its a damn Test

OpenStudy (immanuelv):

switch x and the y then solve for y again

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

can you help me with my whole test???

OpenStudy (immanuelv):

for 5 min... gotta sleep!

OpenStudy (immanuelv):

so what is the question

OpenStudy (anonymous):

there are alot more ..im on 1B

OpenStudy (immanuelv):

lol... but i have a semester exam tomorrow.... so hurry.

OpenStudy (anonymous):

ohhhh.....well im doin a Take home test rite now

OpenStudy (immanuelv):

k

OpenStudy (anonymous):

can you tell me the Point i would be using if i used 2, 1, 0, -1, -2

OpenStudy (anonymous):

and those numbers were x

OpenStudy (immanuelv):

??

OpenStudy (anonymous):

like in the Equation you gave me

OpenStudy (anonymous):

but put those numbers one by one for X and whats the answer everytime

OpenStudy (immanuelv):

for the equation i gave you?

OpenStudy (anonymous):

yea...the inverse

OpenStudy (anonymous):

for thise 1±((√3−x))/((√2))

OpenStudy (anonymous):

can you tell me what the answer is if you plug the 5numbers i told you

OpenStudy (immanuelv):

2= 1.707 and .2928 1= 2 and 0 0= 2.225 and -.2247 -1= 2.414 and -.414 -2 = 2.581 and -.581

OpenStudy (immanuelv):

is that helpful?

OpenStudy (immanuelv):

its 64

OpenStudy (anonymous):

huh?

OpenStudy (immanuelv):

sorry wrong post!

OpenStudy (anonymous):

whats the inverse of 4^x

OpenStudy (immanuelv):

(log x)/ (log 4)

OpenStudy (anonymous):

no its f(x)=4^x...i need inverse of that

OpenStudy (immanuelv):

it is its inverse!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x) = 4^x\]\[f(f^{-1}(x)) = x\implies x = 4^{f^{-1}(x)} \]\[\implies ln\ x = f^{-1}(x)( ln\ 4) \]\[\implies f^{-1}(x) = {ln\ x \over ln\ 4}\]

OpenStudy (immanuelv):

i was right :)